# Exploring Value Biases: How LLMs Deviate Towards the Ideal

### 16 Feb 2024 v1

@@ 4.3 Value-bias in prototypes
These exemplars are evaluated on the three dimensions of averageness, idealness, and prototypicality using the same format used in the human study Bear & Knobe (2017b). Prototypicality is further divided into three entities, which measure the degree to which the given prototype is a “good example”, “paradigmatic example”, or “prototypical example” as done in the human study. The prompt for the five conditions follows the same format across the eight different categories ($C$s).
@@
The prompt above gets the LLM to rate “how average the exemplar is”. Other prompts are formed similarly; the prompt set is given in Appendix 1.5
We run this experiment ten times for a default temperature of 0.8 and report the average results. The average scores from the three prototypicality assessments (“good”, “paradigmatic”, and “prototypical” example) demonstrated satisfactory internal consistency, with a Cronbach’s $\alpha$ of 0.96. Consequently, these scores were combined to form a single, comprehensive prototypicality rating and the aggregate results averaged across exemplars are given in Table 2. The complete set of results for every exemplar is given in the Appendix 1.6.
@@ 5 Discussions
The terms used in this section are not intended to anthropomorphize the LLM; we use terminology like ’the notion of ideal’ for the lack of a better phrase. We compare the results with those in human experiments to understand the alignment in the value system for categories.
@@
Comparing the experiment 4.1 with the same categories done with human subjects by  Bear et al. (2020) in Appendix 1.1, the LLM often gives a ‘strictly ideal’ value when queried for $C_{i}$. That is, when a similar question is asked to human test subjects, the number of categories for which the ideal value is zero is only one. On the other hand, LLM gives zero for $C_{i}$ for 19 categories (nearly half the time). For instance, the human gives the ideal percentage of ‘high school students underage drinking’ as 13.71% while the LLM gives $C_{i}$ as zero for this category.
@@
Figure 4(a) shows the scatter plot between the $\hat{\alpha}$ value for LLMs and humans. We can see that although LLM has a strong value bias based on its implicit value associated with each category, its value system does not correlate with that of humans (Pearson correlation of -0.02). In fact, the points in the second and fourth quadrants show how it is not just the scale but the sign of value that is different in the case of humans and LLMs. This makes the study of value bias in LLM more significant as it might not align with human value systems more often than they align.
@@
Comparing $\hat{\alpha}$ of humans and LLM for Experiment 4.3 shows a higher alignment in the value in Figure 4(b). Here the Pearson correlation of $\hat{\alpha}_{human}$ and $\hat{\alpha}_{LLM}$ is 0.33. Though not fully aligned in many categories, only two have different polarities for $\hat{\alpha}$. This shows the high agreement in valence in these categories. Across the two studies comparing humans and LLM, though the results don’t align, it motivates the need to study value bias in contexts where they get deployed.
Impact of value bias on other applications
@@
We hypothesize that value bias can manifest in many different applications. Consider a simplified version of summarization where the LLM is asked to summarize a given description by picking just one adjective and summarizing it in one sentence. We use prompts similar to Experiment 3 Bear & Knobe (2017b) but with mixed valences within a description. We ask an expert to create these prompts. We anecdotally observe that despite the mixed valences in the description, LLM constantly samples positive valence adjectives to summarize the multi-line description with mixed valence. For this study, we chose only categories with positive valence. Detailed study into different aspects of value bias can help mitigate these risks in real-world applications. Results in Appendix 1.7.
Value bias across the studies
@@
Increasingly, LLMs are being used as agents and decision-making systems deployed in the real world with significant impact Wang et al. (2023). Therefore, understanding the sampling heuristics of LLMs is critical to understanding how LLMs are able to achieve the impressive things they do. They seem to utilize some sort of heuristics to narrow down a vast list of possibilities into a useful high-level plan across a range of domains Romera-Paredes et al. (2023). Our results seem to suggest LLMs, through their token prediction over a large training data, might have acquired some internal notion of value or goal with it Andreas (2022).
@@
The results of experiment 4.1 show that LLMs have an implicit value bias in sampling. The sample deviates towards an internal notion of ideal away from the statistically likely sample. Across multiple scenarios evaluated, LLMs seem to sample a higher value if the task is considered in its value function $V$ as positive ($C_{i}>C_{a}$)and vice versa, sampling a lower value when the task was negative ($C_{i}<C_{a}$) for the LLM.
@@
Figure 4: (Left) The two values are not correlated. The $\hat{\alpha}$ for LLM plotted against the $\hat{\alpha}$ computed from the human evaluation. The two values are correlated. (Right) The $\hat{\alpha}$ for LLM plotted against the $\hat{\alpha}$ computed from the human evaluation for Experiment  4.1.
Psychological studies in humans
indicate that humans sample from a probability distribution that maximizes both frequency and value Bear & Knobe (2017b)Bear et al. (2020). It has been hypothesized such an adaptation could be helpful for decision-making in humans to filter down possibilities efficiently from the vast search space Phillips et al. (2019). Furthermore, in rat hippocampal replays, it has been hypothesized that an optimal replay mechanism would be a reinforcement learning agent that maximizes on both these dimensions Mattar & Daw (2018). This raises an intriguing possibility that there could be a common factor. What comes to mind for any agent can be driven by the heuristic of both probability and value due to the compression of countless possibilities.
@@ 6 Conclusion
LLMs are being used in multiple domains across various applications. With its increasing impact, understanding the mechanism by which LLM samples responses is critical. In this paper, we study the sampling of LLM in three different settings. We observe that the sample has a shift from what is statistically likely towards an ideal value under some value system. This causes the LLM to show a value bias. We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero-shot). We compare the value bias of LLMs to humans and show that the two might not always align. As a final remark, we would like to emphasize that we do not intend to contribute to “humanizing” AI/ML/LLMs in the way we use terminology or models. Our contribution is intended to draw parallels in behaviour and perform evaluations, as our findings will have an impact on downstream tasks.

### 21 Feb 2024 v2

@@ 4.3 Value Bias in Prototypes
These exemplars are evaluated on the three dimensions of averageness, idealness, and prototypicality using the same format used in the human study (Bear & Knobe, 2017b). Prototypicality is further divided into three entities, which measure the degree to which the given prototype is a “good example”, “paradigmatic example”, or “prototypical example” as done in the human study. The prompt for the five conditions follows the same format across the eight different categories ($C$s).
@@
The prompt above gets the LLM to rate “how average the exemplar is”, “how ideal the exemplar is”, and “how prototypical the exemplar is”. The complete set of exemplars is given in the Appendix 1.5. A prompt is appended to each question to get the rating. A sample for average prompt is shown below:
We run this experiment ten times for a default temperature of 0.8 and report the average results. The average scores from the three prototypicality assessments (“good”, “paradigmatic”, and “prototypical” example) demonstrated satisfactory internal consistency, with a Cronbach’s $\alpha$ of 0.96. Consequently, these scores were combined to form a single, comprehensive prototypicality rating and the aggregate results averaged across exemplars are given in Table 2. The complete set of results for every exemplar is given in Appendix 1.6.
@@ 5 Discussions
The terms used in this section (and the paper) are not intended to anthropomorphize the LLM; we use terminology like ‘the notion of ideal’ for the lack of a better phrase. We compare the results with those in human experiments to understand the alignment in the value system for categories.
@@
Comparing experiment 4.1 with the same categories done with human subjects by Bear et al. (2020) in Appendix 1.1, the LLM often gives a ‘strictly ideal’ value when queried for $C_{i}$. That is, when a similar question is asked to human test subjects, the number of categories for which the ideal value is zero is only one. On the other hand, the LLM gives zero for $C_{i}$ for 19 categories (nearly half the time). For instance, the human gives the ideal percentage of ‘high school students underage drinking’ as 13.71% while the LLM gives $C_{i}$ as zero for this category.
@@
Figure 4(a) shows the scatter plot between the $\hat{\alpha}$ value for LLMs and humans. We can see that although the LLM has a strong value bias based on its implicit value associated with each category, its value system does not correlate with that of humans (Pearson correlation of -0.02). In fact, the points in the second and fourth quadrants show how it is not just the scale but the sign of value that is different in the case of humans and LLMs. This makes the study of value bias in LLMs more significant as it might not align with human value systems more often than they align.
@@
Comparing $\hat{\alpha}$ of humans and the LLM for experiment 4.3 shows a higher alignment in the value in Figure 4(b). Here the Pearson correlation of $\hat{\alpha}_{human}$ and $\hat{\alpha}_{LLM}$ is 0.33. Though not fully aligned in many categories, only two have different polarities for $\hat{\alpha}$. This shows the high agreement in valence in these categories. The two studies comparing humans and LLMs, where the results do not align, motivate the need to study value bias in contexts where LLMs get deployed.
Impact of value bias on other applications.
@@
We hypothesize that value bias can manifest in many different applications. Consider a simplified version of summarization where the LLM is asked to summarize a given description by picking just one adjective and summarizing it in one sentence. We use prompts similar to Experiment 3 (Bear & Knobe, 2017b) but with mixed valences within a description. We ask an expert to create these prompts. We anecdotally observe that despite the mixed valences in the description, the LLM constantly samples positive valence adjectives to summarize the multi-line description with mixed valence. For this study, we chose only categories with positive valence. A further detailed study of the different aspects of value bias can help mitigate these risks in real-world applications. Results in Appendix 1.7.
Value bias across the studies.
@@
Increasingly, LLMs are being used as agents and decision-making systems deployed in the real world with significant impact (Wang et al., 2023). Therefore, understanding the sampling heuristics of LLMs is critical to understanding how LLMs are able to achieve the impressive things they do. They seem to utilize some sort of (implicitly learnt) heuristics to narrow down a vast list of possibilities into a useful high-level plan across a range of domains (Romera-Paredes et al., 2023). Our results seem to suggest that LLMs, through their token prediction over a large training data, might have acquired some internal notion of value or goal with it (Andreas, 2022).
@@
The results of experiment 4.1 show that LLMs have an implicit value bias in sampling. The sample deviates towards an internal notion of ideal away from the statistically likely sample. Across multiple scenarios evaluated, LLMs seem to sample a higher value if the task is considered by its value function $V$ as positive ($C_{i}>C_{a}$) and, vice versa, sampling a lower value when the task is negative ($C_{i}<C_{a}$).
@@
Figure 4: (Left) shows the comparison of $\hat{\alpha}$ for LLM and human on Experiment  4.1. The two values are not correlated. (Right) The $\hat{\alpha}$ for the LLM plotted against the $\hat{\alpha}$ computed from the human evaluation for Experiment 4.3.
Psychological studies in humans.
These indicate that humans sample from a probability distribution that maximizes both frequency and value (Bear & Knobe, 2017b; Bear et al., 2020). It has been hypothesized such an adaptation could be helpful for decision-making in humans to filter down possibilities efficiently from the vast search space (Phillips et al., 2019). Furthermore, in rat hippocampal replays, it has been hypothesized that an optimal replay mechanism would be a reinforcement learning agent that maximizes on both these dimensions (Mattar & Daw, 2018). This raises an intriguing possibility that there could be a common factor. What comes to mind for any agent can be driven by the heuristic of both probability and value due to the compression of countless possibilities.
@@ 6 Conclusion
LLMs are being used in multiple domains across various applications. With its increasing impact, understanding the mechanism by which am LLM samples responses is critical. In this paper, we study the sampling of LLMs in three different settings. We observe that the sample has a shift from what is statistically likely towards an ideal value under some value system. This causes the LLM to show a value bias. We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero-shot). We compare the value bias of LLMs to humans and show that the two might not always align. As a final remark, we would like to emphasize that we do not intend to contribute to “humanizing” AI/ML/LLMs in the way we use terminology or models. Our contribution is intended to draw parallels in behaviour and perform evaluations, as our findings will have an impact on downstream tasks.